OpenStudy (anonymous):

30 mi / h = ____ ft / min

OpenStudy (jack1):

\[\frac{ 30 miles }{ 1hr } \times \frac{ 1 hr }{ 60 \min } \times \frac{ 5280 ft }{ 1 mile } = answer\]

OpenStudy (jack1):

\[ = \frac{ 30\times1\times5280 }{ 1\times60\times1 }\]